OpenStudy (sonyalee77):

What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6? A. y=-3/2x+8 B. y=3/2x-4 C.y=1/2x D. y=2/3x-4

OpenStudy (anonymous):

3x - 2y = -6 has gradient -3/-2 = 3/2 so equation of line is also 3/2 use the point slope formula to find the equation

OpenStudy (anonymous):

y - y1 = m (x - x1)

OpenStudy (anonymous):

* the gradient of the equation of line is also 3/2

OpenStudy (anonymous):

the answer should be B